OpenStudy (anonymous):

what is the length (magnitude) of the vector (-4 2) A. 5 B. 3 C. root 2 D. root 20

OpenStudy (danjs):

The length of the hypotenuse of a triangle, with side lengths of 4 and 2

OpenStudy (danjs):

\[\left| \left| <-4,2> \right| \right| = \sqrt{(-4)^2+2^2}\]

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

its precal @DanJS

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

root 20? @DanJS

OpenStudy (triciaal):

yes